OpenStudy (anonymous):

Question? 18. If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 6. A. -30 B. -32 C. -34 D. -36

OpenStudy (anonymous):

x=ky y=25, k = 2/5 x=2/5 * 25 x= 10

OpenStudy (anonymous):

sorry but i dont understand the question

OpenStudy (anonymous):

oh its directly so its y= kx

OpenStudy (anonymous):

15=-3k k=-5 y=-5x can you finish?

OpenStudy (anonymous):

forget about the 45 y varies directly so its Y=K*X

OpenStudy (anonymous):

\[y = k*x\] \[y = 15\] \[x = -3\] \[15 = -3*k\] solve for k

OpenStudy (anonymous):

wrong, why do you multiply?

OpenStudy (anonymous):

i will do it step by step \[y=k*x\] divide by x \[\frac{y}{x}=k\] \[\frac{15}{-3}=k\] \[-5=k\] plug it in \[y=-5*x\] \[y=-5*6\] \[y=-30\] thats why you should do the problems yourself and not trust every answer

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

correct